A plumb bob of mass 1 kg is hung from the ceiling of a train compartment. The train moves with an acceleration of g/2 up on an inclined plane. If the angle of inclination is 30 degrees. Find the angle made by the string with the normal to the ceiling. Also find the tension in the string. (g=10 m/s.s)

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

component of force on bob perpendicular to normal= (g/2)+(gsin30)=g,component parallel=gcos30=g√3/2,angle=arctan(g/(g√3/2))=arctan(2/√3),tension=sqrt(g^2 +(3/4)g^2)=(g/2)√7 N